OpenStudy (anonymous):

Evaluate 2x^2y for x = 2 1/2 and y = -3 3/5

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

i plugged in 2(2 1/2) ^2 (-3 3/5)

OpenStudy (anonymous):

\[2(2\frac{ 1 }{ 2 })^2 (-3\frac{ 3 }{ 5})\]

OpenStudy (jhannybean):

I think what will make it a little easier to evaluate would be if you changed the mixed integer into a fraction. \[2\frac{1}{2} \iff \frac{(2 \cdot 2)+1}{2}\]\[-3\frac{3}{5} \iff \frac{(-3 \cdot 5) +3}{5}\]

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

and from there do i just multiply straight across?

OpenStudy (jhannybean):

Yup, you input the fraction in place of x and y and simplify! :)

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

okay so after that i got 5/2 and then -12/5 is that right?

OpenStudy (jhannybean):

Lets check it out. I havent solved it yet! SO we'll see.

OpenStudy (anonymous):

okay! im solving it right now

OpenStudy (anonymous):

i figured it out! thanks

OpenStudy (jhannybean):

\[x = \frac{ 5}{2} ~,~ y = -\frac{12}{5}\]Now let's input these into our formula: \(2x^2y\) When I put these values in, I got:\[2\left(\frac{5}{2}\right)^2\left(\frac{-12}{5}\right)\]\[2\left(\frac{25}{4}\cdot \frac{-12}{5}\right)\]\[2(-15) = \boxed{-30}\]

OpenStudy (anonymous):

Hmm. i got the answer as -45

OpenStudy (jhannybean):

can you write out your steps? Maybe we can compare and see which one of is wrong :o

OpenStudy (anonymous):

I did 2(5^2/2^2) and then -18/5 because i multiped 2*2 then added 1 which gave me 5 so i then put 2(25/4) (-18/5) and then put the 2 into a fraction so 2/1 (25/4) (-18/5) and then i crossed out numbers so i did (2*1/1) (25/2*2) and then it would be 25/2 (-18/5) Cross out numbers again (5*5)/2) (-1*18/5*1) (5/2) (-18/1) = -90/2 = -45

OpenStudy (anonymous):

sorry if you cant understand

OpenStudy (jhannybean):

Where did the \(-\frac{18}{5}\) come from?

OpenStudy (anonymous):

i multiplied -3*5 and added 3

OpenStudy (anonymous):

thank you for helping me! I have to go to school now, the answer sheet said it was -45 so maybe i did something wrong : )

OpenStudy (jhannybean):

Well... \((-3\cdot 5) +3 = -15 + 3 = -12...\)

OpenStudy (anonymous):

yeah i moved the negative sign tho